### v1

@@ 3.3 Combining the iterative and asymptotic methods
For deciding the regions where each of the two methods is used in the final algorithm we follow a similar strategy as for the Gauss-Jacobi case.
@@
Fig. 7 shows the regions where the iterative method is more accurate than the asymptotic method and vice-versa. As measure of the error, we take the maximum relative error for the scaled weights (with unitary normalization); the other two error measures considered before in the Gauss-Jacobi case (average relative error and relative total error) give similar pictures but favoring slightly more the asymptotic approach. We give preference to the iterative method because it is a faster and so we consider the error criteria favoring this method.
Figure 7: Comparison of the maximum relative error of the iterative and the asymptotic method for computing Gauss-Laguerre quadratures. At the red points the iterative method is more accurate that the asymptotic method, while the opposite is true both for the green and the blue points; at the blue points the accuracy of the iterative method is ten times worse than that of the asymptotic method. The separation lines between the methods used in the combined algorithm are also shown.
@@
For the rest of parameters we use the iterative method except $n\leq 5$, where we use the GW method. We do not consider the iterative method for $n<5$ because the method was not originally designed for such small values of $n$ (see [10, section 4.2 and 4.3 ] for further details). A version for $n<5$ is possible, but GW does the job perfectly for any $\alpha$ (no scaled weights are produced, though).
@@
As commented in the Gauss-Jacobi case, the accuracy of the Bessel-type asymptotic expansions is limited by the accuracy in the computation of Bessel functions with the intrinsic Matlab procedure. For the Gauss-Laguerre case, this limitation is of higher importance because it affects the smallest positive nodes, and therefore the most significant weights for small $\alpha$. In order to improve the accuracy of the algorithm we use our own algorithms (included in the package) for the computation of Bessel functions, which prove to be more accurate that the Matlab intrinsic implementation, at least for small orders; we use our expansions for $1<\alpha\leq 5$ and use the intrinsic Matlab codes for $\alpha>5$.
@@ 3.4 Testing the combined algorithm
4 Gauss-Hermite quadrature
@@
The nodes $x_{i}$ of $n$-point Gauss-Hermite quadrature are the roots of $H_{n}(x)$, which are symmetric around the origin, and the weights $w_{i}$ can be written as (see Table 1):
@@ 4.1 Iterative method
The iterative method for Gauss-Hermite is a simple, accurate and efficient method of computation. The nodes are computed with full double accuracy, while some mild degradation is observed in the computation of the weights as the degree increases. For computing scaled weights for large $n$, the asymptotic methods produce higher accuracy. However, as we discuss later, the iterative method for Gauss-Hermite also has a number of advantages with respect to asymptotics.
@@ 4.2 Asymptotic methods
We describe here some numerical aspects of the asymptotic computation which were not described in [9] and which are important for an acurate computation, particularly of the weights.
@@ 4.2.1 Computation of the Airy-type expansion
where in the last equation we have used the relation between Hermite polynomials and parabolic cylinder functions [22, 12.7.2] $y(x)=2^{n/2}U(-n-\frac{1}{2},\sqrt{2}x)$. As we did earlier in the Laguerre case, we need to give a brief description regarding the asymptotic computation of the derivative $y^{\prime}(x_{i})$. For this purpose, we use the Airy-type expansion for Parabolic Cylinder functions (see [22], section 12.10(vii)).
@@
where $\frac{1}{3}(-\zeta)^{3/2}=\frac{\displaystyle{1}}{\displaystyle{2}}\left(\arccos t-t\sqrt{1-t^{2}}\right)$ in the region where the nodes lie ($|t|<1$) and $g(\mu)$, $\phi(\zeta)$ are as defined in [22, section 12.20]. The coefficients admit the expansions
@@
With respect to the computation of the Airy functions, a similar procedure as described for the Gauss-Laguerre case can be used here. Indeed, the dominant term in the asymptotic approximation for the nodes is determined by setting to zero the dominant contribution in the expansion for the Hermite polynomial [22, 12.10.35]. If $\zeta_{i}$ is the value of $\zeta$ corresponding to the node $x_{i}$, we have $\mu^{4/3}\zeta_{n-i+1}=a_{i}+\delta_{i}$, where $\delta_{i}$ can be computed accurately and $\delta_{i}/a_{i}<<1$. We can compute the Airy functions in (31) using Taylor series. For Gauss-Hermite quadrature the Airy-type expansion is only used for the last $20$ weights and we store the precomputed values of $a_{i}$ and ${\rm Ai}^{\prime}(ai)$, $i=1,\ldots 20$.
@@ 4.3 Comparing the iterative and asymptotic methods
The accuracy of the iterative and asymptotic methods is compared in Table 2. The nodes are computed with slightly more accuracy for the iterative method, which produces full double precision accuracy for any degree. Contrarily, the scaled weights are more accurately computed with the asymptotic method when $n\geq 150$ than with the iterative method. However, the maximum relative error for the unscaled weights which are computable in double precision (limited by underflow) is very similar for the iterative and asymptotic method. The main source of error for the unscaled nodes is the exponential factor $e^{-x_{i}^{2}}$, which is slightly more accurately computed with the iterative method.
@@ 5 Further applications
In this section, we discuss how to compute the Gauss-Lobatto and Gauss-Radau quadratures, as well as the barycentric weights associated to the different sets of classical nodes (including Gauss-Lobatto and Gauss-Radau variants).
5.1 Gauss-Radau and Gauss-Lobatto quadratures
In some applications, for instance for solving boundary problems, it can be interesting to consider rules with the maximal degree of polynomial exactness which include one or two of the endpoints of the interval of integration as nodes of integration. When only one of the endponts is considered the rule is called of Gauss-Radau type, and when the two endpoints are used of Gauss-Lobatto type.
Let us first consider a generic Gauss-Radau case for a weight function $w(x)$ in an interval $[a,b]$ with $x=a$ as a node:
@@ 5.1 Gauss-Radau and Gauss-Lobatto quadratures
For the Gauss-Radau case with $x=b$ as a node we proceed similarly. For the Gauss-Lobatto case, with both $x=a$ and $x_{b}$ as nodes, the internal nodes are the zeros of the polynomial $p_{n}^{L}(x)$ orthogonal with respect $w^{L}(x)=(x-a)(b-x)w(x)$, and the relation between weights follow similarly: $w_{i}=w_{i}^{L}/[(x_{i}-a)(b-x_{i})]$.
As for the boundary node(s), and considering again the Gauss-Radau case with $x=a$ as a node, because $p^{R}_{n}(x_{i})=0,\,i=1..n$ and $Q_{n}$ has degree of exactness $2n$
@@
For the Gauss-Lobatto case we would consider the integration of $(x-a)p_{n}^{L}(x)$ and $(b-x)p_{n}^{L}(x)$ for determining the boundary nodes.
For the cases of classical Gaussian rules the polynomials $p_{n}^{R}(x)$ are classical orthogonal polynomials themselves, and the boundary nodes can be computed explicitly. Take for instance the case of Gauss-Radau-Laguerre quadrature, which is the Radau case considered before with $a=0$, $b=+\infty$, $w(x)=x^{-\alpha}e^{-x}$. The Laguerre weight is $w(x)=x^{\alpha}e^{-x}$, and in this case $p_{n}^{R}(x)=L_{n}^{(\alpha+1)}(x)$ is the orthogonal polynomial associated to the weight $w^{R}(x)=xw(x)$. To compute the Gauss-Radau-Laguerre quadrature, we first compute the nodes $x_{i}^{R}$ and weights $w_{i}^{R}$ of Gauss-Laguerre quadrature with parameter $\alpha+1$; then the internal nodes of the Gauss-Radau-Laguerre quadrature with parameter $\alpha$ are $x_{i}=x_{i}^{R}$, $i=1,\ldots n$, and the internal weights are $w_{i}=w_{i}^{R}/x_{i}$, $i=1,\ldots n$. The boundary weight, considering (36) is
@@
and because $L_{n}^{(\alpha+1)}(x)=\sum_{j=0}^{n}L_{j}^{(\alpha)}(x)$ [16, 18.18.37], using the orthogonality of Laguerre polynomials and the fact that $L_{n}^{(\alpha)}(0)=(\alpha+1)_{n}/n!$ we have
@@
For the rest of Gauss-Radau and Gauss-Lobatto quadratures the computation of the boundary weights follow similar ideas, and the main task is to compute, as in (37), the integral of an orthogonal polynomial of the same family associated to the weight function, but with one displaced parameter. For details, we mention [6] for Gauss-Radau formulae and [7] for Gauss-Lobatto, and we only summarize the results:
  1. 1.
    Gauss-Laguerre-Radau. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss-Laguerre formula with parameter $\alpha+1$, then the quadrature rule
@@
  2. 2.
    Gauss-Jacobi-Radau with x=1x=-1 as a node. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss-Jacobi formula with parameters $\alpha$ and $\beta+1$, then the quadrature rule
@@
  3. 3.
    Gauss-Jacobi-Radau with $x=+1$ as a node. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss-Jacobi formula with parameters $\alpha+1$ and $\beta$, then the quadrature rule
@@
  4. 4.
    Gauss-Jacobi-Lobatto. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss-Jacobi formula with parameters $\alpha+1$ and $\beta+1$, then the quadrature rule
@@ 5.2 Barycentric interpolation
In our algorithms, we compute barycentric nodes and weights for the classical Gaussian quadratures, including also the Gauss-Radau and Gauss-Lobatto variants. Just by comparing the expression of the Gaussian weights with (44) we can relate the barycentric weights with the Gaussian weights. $\Phi^{\prime}(x_{i})$ in (44) is the derivative of the monic orthogonal polynomial at the node $x_{i}$, and this same quantity appears in the expression for the Gauss weights (squared). Neglecting constant multiplicative factors (not depending on the nodes $x_{i}$), the relation is thus very simple. For the Hermite case we can write $u_{i}=(-1)^{i}\sqrt{w_{i}}$ where the $(-1)^{i}$ is due to the fact that, being all the zeros of $\Phi(x)$ simple, the sign of the derivative is opposite in consecutive zeros; similarly, for the Laguerre case $u_{i}=(-1)^{i}\sqrt{x_{i}w_{i}}$ and for the Jacobi case $u_{i}=(-1)^{i}\sqrt{(1-x_{i}^{2})w_{i}}$.
It is also possible to relate the Gauss and barycentric weights for Gauss-Radau and Gauss-Laguerre quadratures (see [24]). We summarize these results in Table 2.
@@
Table 3: Relation of the barycentric weights with the Gauss (first row), Gauss-Radau (second and third rows) and Gauss-Lobatto (fourth row) weights.
@@ 6 Brief description of the Matlab and Maple algorithms
We provide Matlab implementations of the three classical Gaussian quadratures together with arbitrary accuracy Maple algorithms for the symmetric cases (Hermite, Gegenbauer). Gauss-Radau and Gauss-Lobatto quadratures are also implemented, together with the computation of barycentric weights. The codes are available at github.com/jjavier-segura/GaussQuadrature.
The following programs are provided:
6.1 MATLAB algorithms
  1. 1.
    Gauss-Jacobi quadratures
    1. (a)
      GJ.m: computation of Gauss-Jacobi quadrature (uses legen.m).
    2. (b)
      legen.m: fast computation of Gauss-Legendre quadrature.
    3. (c)
      rlJ.m: Gauss-Radau-Jacobi and Gauss-Lobatto-Jacobi quadratures and barycentric weights (uses GJ.m).
  2. 2.
    Gauss-Laguerre quadratures
    1. (a)
      GL.m: computation of Gauss-Laguerre quadrature (uses bessJY.m).
    2. (b)
      rL.m: Gauss-Radau-Laguerre and barycentric weights (uses GL.m).
  3. 3.
    Gauss-Hermite quadrature
    1. (a)
      GHi.m: iterative computation of Gauss-Hermite quadrature.
    2. (b)
      GHa.m: asymptotic computation of Gauss-Hermite quadrature.
    3. (c)
      bH.m: computation of barycentric weights.
6.2 MAPLE algorithms
  1. 1.
    Gauss-Gegenbauer quadrature: gegenbauer.mpl (Maple algorithm) and gegenbauer.mws (worksheet illustrating the use of gegenbauer.mpl).
  2. 2.
    Gauss-Hermite quadrature: hermite.mpl and hermite.mws (Maple algorithm and illustrative Maple worksheet).

### v2

@@ 3.3 Combining the iterative and asymptotic methods
For deciding the regions where each of the two methods is used in the final algorithm we follow a similar strategy as for the Gauss–Jacobi case.
@@
Fig. 7 shows the regions where the iterative method is more accurate than the asymptotic method and vice-versa. As measure of the error, we take the maximum relative error for the scaled weights (with unitary normalization); the other two error measures considered before in the Gauss–Jacobi case (average relative error and relative total error) give similar pictures but favoring slightly more the asymptotic approach. We give preference to the iterative method because it is a faster and so we consider the error criteria favoring this method.
Figure 7: Comparison of the maximum relative error of the iterative and the asymptotic method for computing Gauss–Laguerre quadratures. At the red points the iterative method is more accurate that the asymptotic method, while the opposite is true both for the green and the blue points; at the blue points the accuracy of the iterative method is ten times worse than that of the asymptotic method. The separation lines between the methods used in the combined algorithm are also shown.
@@
For the rest of parameters we use the iterative method except $n\leq 5$, where we use the GW method. We do not consider the iterative method for $n<5$ because the method was not originally designed for such small values of $n$ (see [14, section 4.2 and 4.3 ] for further details). A version for $n<5$ is possible, but GW does the job perfectly for any $\alpha$ (no scaled weights are produced, though).
@@
As commented in the Gauss–Jacobi case, the accuracy of the Bessel-type asymptotic expansions is limited by the accuracy in the computation of Bessel functions with the intrinsic MATLAB procedure. For the Gauss–Laguerre case, this limitation is of higher importance because it affects the smallest positive nodes, and therefore the most significant weights for small $\alpha$. In order to improve the accuracy of the algorithm we use our own algorithms (included in the package) for the computation of Bessel functions, which prove to be more accurate that the MATLAB intrinsic implementation, at least for small orders; we use our expansions for $1<\alpha\leq 5$ and use the intrinsic MATLAB codes for $\alpha>5$.
@@ 3.4 Testing the combined algorithm
4 Gauss–Hermite quadrature
@@
The nodes $x_{i}$ of $n$-point Gauss–Hermite quadrature are the roots of $H_{n}(x)$, which are symmetric around the origin, and the weights $w_{i}$ can be written as (see Table 1):
@@ 4.1 Iterative method
The iterative method for Gauss–Hermite is a simple, accurate and efficient method of computation. The nodes are computed with full double accuracy, while some mild degradation is observed in the computation of the weights as the degree increases. For computing scaled weights for large $n$, the asymptotic methods produce higher accuracy. However, as we discuss later, the iterative method for Gauss–Hermite also has a number of advantages with respect to asymptotics.
@@ 4.2 Asymptotic methods
We describe here some numerical aspects of the asymptotic computation which were not described in [13] and which are important for an accurate computation, particularly of the weights.
@@ 4.2.1 Computation of the Airy-type expansion
where in the last equation we have used the relation between Hermite polynomials and parabolic cylinder functions [30, 12.7.2] $y(x)=2^{n/2}U(-n-\frac{1}{2},\sqrt{2}x)$. As we did earlier in the Laguerre case, we need to give a brief description regarding the asymptotic computation of the derivative $y^{\prime}(x_{i})$. For this purpose, we use the Airy-type expansion for Parabolic Cylinder functions (see [30], section 12.10(vii)).
@@
where $\frac{1}{3}(-\zeta)^{3/2}=\frac{\displaystyle{1}}{\displaystyle{2}}\left(\arccos t-t\sqrt{1-t^{2}}\right)$ in the region where the nodes lie ($|t|<1$) and $g(\mu)$, $\phi(\zeta)$ are as defined in [30, section 12.20]. The coefficients admit the expansions
@@
With respect to the computation of the Airy functions, a similar procedure as described for the Gauss–Laguerre case can be used here. Indeed, the dominant term in the asymptotic approximation for the nodes is determined by setting to zero the dominant contribution in the expansion for the Hermite polynomial [30, 12.10.35]. If $\zeta_{i}$ is the value of $\zeta$ corresponding to the node $x_{i}$, we have $\mu^{4/3}\zeta_{n-i+1}=a_{i}+\delta_{i}$, where $\delta_{i}$ can be computed accurately and $\delta_{i}/a_{i}<<1$. We can compute the Airy functions in (32) using Taylor series. For Gauss–Hermite quadrature the Airy-type expansion is only used for the last $20$ weights and we store the precomputed values of $a_{i}$ and ${\rm Ai}^{\prime}(ai)$, $i=1,\ldots 20$.
@@ 4.3 Comparing the iterative and asymptotic methods
The accuracy of the iterative and asymptotic methods is compared in Table 2. The nodes are computed with slightly more accuracy for the iterative method, which produces full double precision accuracy for any degree. Contrarily, the scaled weights are more accurately computed with the asymptotic method when $n\geq 150$ than with the iterative method. However, the maximum relative error for the unscaled weights which are computable in double precision (limited by underflow) is very similar for the iterative and asymptotic method. The main source of error for the unscaled weights is the exponential factor $e^{-x_{i}^{2}}$, which is slightly more accurately computed with the iterative method.
@@ 5 Further applications
In this section, we discuss how to compute the Gauss–Lobatto and Gauss–Radau quadratures, as well as the barycentric weights associated to the different sets of classical nodes (including Gauss–Lobatto and Gauss–Radau variants).
5.1 Gauss–Radau and Gauss–Lobatto quadratures
In some applications, for instance for solving boundary problems, it can be interesting to consider rules with the maximal degree of polynomial exactness which include one or two of the endpoints of the interval of integration as nodes of integration. When only one of the endpoints is considered the rule is called of Gauss–Radau type, and when the two endpoints are used of Gauss–Lobatto type.
Let us first consider a generic Gauss–Radau case for a weight function $w(x)$ in an interval $[a,b]$ with $x=a$ as a node:
@@ 5.1 Gauss–Radau and Gauss–Lobatto quadratures
For the Gauss–Radau case with $x=b$ as a node we proceed similarly. For the Gauss–Lobatto case, with both $x=a$ and $x_{b}$ as nodes, the internal nodes are the zeros of the polynomial $p_{n}^{L}(x)$ orthogonal with respect $w^{L}(x)=(x-a)(b-x)w(x)$, and the relation between weights follow similarly: $w_{i}=w_{i}^{L}/[(x_{i}-a)(b-x_{i})]$.
As for the boundary node(s), and considering again the Gauss–Radau case with $x=a$ as a node, because $p^{R}_{n}(x_{i})=0,\,i=1..n$ and $Q_{n}$ has degree of exactness $2n$
@@
For the Gauss–Lobatto case we would consider the integration of $(x-a)p_{n}^{L}(x)$ and $(b-x)p_{n}^{L}(x)$ for determining the boundary nodes.
For the cases of classical Gaussian rules the polynomials $p_{n}^{R}(x)$ are classical orthogonal polynomials themselves, and the boundary nodes can be computed explicitly. Take for instance the case of Gauss–Radau–Laguerre quadrature, which is the Radau case considered before with $a=0$, $b=+\infty$, $w(x)=x^{-\alpha}e^{-x}$. The Laguerre weight is $w(x)=x^{\alpha}e^{-x}$, and in this case $p_{n}^{R}(x)=L_{n}^{(\alpha+1)}(x)$ is the orthogonal polynomial associated to the weight $w^{R}(x)=xw(x)$. To compute the Gauss–Radau–Laguerre quadrature, we first compute the nodes $x_{i}^{R}$ and weights $w_{i}^{R}$ of Gauss–Laguerre quadrature with parameter $\alpha+1$; then the internal nodes of the Gauss–Radau–Laguerre quadrature with parameter $\alpha$ are $x_{i}=x_{i}^{R}$, $i=1,\ldots n$, and the internal weights are $w_{i}=w_{i}^{R}/x_{i}$, $i=1,\ldots n$. The boundary weight, considering (37) is
@@
and because $L_{n}^{(\alpha+1)}(x)=\sum_{j=0}^{n}L_{j}^{(\alpha)}(x)$ [22, 18.18.37], using the orthogonality of Laguerre polynomials and the fact that $L_{n}^{(\alpha)}(0)=(\alpha+1)_{n}/n!$ we have
@@
For the rest of Gauss–Radau and Gauss–Lobatto quadratures the computation of the boundary weights follow similar ideas, and the main task is to compute, as in (38), the integral of an orthogonal polynomial of the same family associated to the weight function, but with one displaced parameter. For details, we mention [8] for Gauss–Radau formulae and [9] for Gauss–Lobatto, and we only summarize the results:
1. Gauss–Laguerre–Radau. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss–Laguerre formula with parameter $\alpha+1$, then the quadrature rule
@@
2. Gauss–Jacobi–Radau with x=1x=-1 as a node. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss–Jacobi formula with parameters $\alpha$ and $\beta+1$, then the quadrature rule
@@
3. Gauss–Jacobi–Radau with $x=+1$ as a node. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss–Jacobi formula with parameters $\alpha+1$ and $\beta$, then the quadrature rule
@@
4. Gauss–Jacobi–Lobatto. Let $x_{i}^{R}$ and $w_{i}^{R}$ be the nodes and weights for the $n$-point Gauss–Jacobi formula with parameters $\alpha+1$ and $\beta+1$, then the quadrature rule
@@ 5.2 Barycentric interpolation
In our algorithms, we compute barycentric nodes and weights for the classical Gaussian quadratures, including also the Gauss–Radau and Gauss–Lobatto variants. Just by comparing the expression of the Gaussian weights with (45) we can relate the barycentric weights with the Gaussian weights. $\Phi^{\prime}(x_{i})$ in (45) is the derivative of the monic orthogonal polynomial at the node $x_{i}$, and this same quantity appears in the expression for the Gauss weights (squared). Neglecting constant multiplicative factors (not depending on the nodes $x_{i}$), the relation is thus very simple. For the Hermite case we can write $u_{i}=(-1)^{i}\sqrt{w_{i}}$ where the $(-1)^{i}$ is due to the fact that, being all the zeros of $\Phi(x)$ simple, the sign of the derivative is opposite in consecutive zeros; similarly, for the Laguerre case $u_{i}=(-1)^{i}\sqrt{x_{i}w_{i}}$ and for the Jacobi case $u_{i}=(-1)^{i}\sqrt{(1-x_{i}^{2})w_{i}}$.
It is also possible to relate the Gauss and barycentric weights for Gauss–Radau and Gauss–Laguerre quadratures (see [33]). We summarize these results in Table 3.
@@
Table 3: Relation of the barycentric weights with the Gauss (first row), Gauss–Radau (second and third rows) and Gauss–Lobatto (fourth row) weights.
@@ 6 Brief description of the MATLAB and Maple algorithms
[] Matlab algorithms     Gauss-Jacobi quadrature     function GJ: Computation of Gauss-Jacobi quadrature     function legen: Computation of Gauss-Legendre quadrature     function rlJ: Computation of Gauss-Radau-Jacobi and Gauss-Lobatto-Jacobi quadratures and barycentric weights     Gauss-Laguerre quadrature     function GL: Computation of Gauss-Laguerre quadrature     function rL: Computation of Gauss-Radau-Laguerre quadrature and barycentric weights     Gauss-Hermite quadrature     function GHa: Asymptotic computation of Gauss-Hermite quadrature     function GHi: Iterative computation of Gauss-Hermite quadrature     function bH: Computation of barycentric weights     Examples     Maple algorithms     Gauss-Gegenbauer quadrature     gegenbauer.mpl Maple algorithm     gegenbauer.mws illustrative Maple worksheet     Gauss-Hermite quadrature     hermite.mpl Maple algorithm     hermite.mws illustrative Maple worksheet
Figure 9: Overview of the software provided for Gauss quadratures
The codes are available at
The original release (v1.0.0), publicly available at this site, contains the versions used in the present paper.
The different algorithms are documented at the website, including details on the input/output, modes of computation, and error handling; this information is also provided in the comment lines in the headings of the codes. Additionally, examples showing the use of the MATLAB algorithms are provided in the MATLAB folder. Maple worksheets are included in the corresponding folder, demonstrating the use of the arbitrary-precision algorithms for the symmetric cases.
A description of the components of the software library is shown in Figure 9.
Acknowledgments
The authors gratefully acknowledge the two anonymous reviewers for their valuable comments and suggestions.